OpenStudy (anonymous):

Missed a letter from last question, sorry- solve for A: A^2+4a=45

OpenStudy (anonymous):

\[a = - \sqrt{45 - 4a}\] \[a =+ \sqrt{45 - 4a}\]

OpenStudy (anonymous):

a^2=4a-45=0 (a+9)(a-5) a=-9,5

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

ur welcome medal please